OpenStudy (shadowcain):

Which equation illustrates the division property of equality? A. d x 4 = 32 d x 4 / 4 = 32 / 4 d = 8 B. d - 4 = 32 d - 4 + 4 = 32 + 4 d = 36 C. d / 4 = 32 d / 4 x 4 = 32 x 4 d = 128 D. d + 4 = 32 d + 4 - 4 = 32 - 4 d = 28

OpenStudy (shadowcain):

I think it's C because 128 / 4 = 32

jimthompson5910 (jim_thompson5910):

division property of equality if a = b, then a/c = b/c where c is nonzero

jimthompson5910 (jim_thompson5910):

in more common language, you might say "divide both sides by ______" (replace the blank with a number or expression)

OpenStudy (shadowcain):

128 / 4 x 4 = 128 32 x 4 = 128

jimthompson5910 (jim_thompson5910):

if we start with d x 4 = 32 we divide both sides by 4 to isolate d this is using the division property of equality

OpenStudy (shadowcain):

They all come out as equal but C is the only one that is dividing. A is multiplication, B. is subtraction and D is adding.

jimthompson5910 (jim_thompson5910):

what you are showing (in choice C) is the multiplicative property of equality

OpenStudy (shadowcain):

Well then I'm totally confused

jimthompson5910 (jim_thompson5910):

you don't start with dividing division is the second step

OpenStudy (shadowcain):

So the division property of equality starts with the opposite, which would be multiplication.

jimthompson5910 (jim_thompson5910):

yeah because the goal is to undo the multiplication

jimthompson5910 (jim_thompson5910):

start with this